### v2

# ToddlerBot: Open-Source ML-Compatible Humanoid Platform for Loco-Manipulation

Abstract
Learning-based robotics research driven by data demands a new approach to robot hardware design—one that serves as both a platform for policy execution and a tool for embodied data collection to train policies. We introduce ToddlerBot, a low-cost, open-source humanoid robot platform designed for scalable policy learning and research in robotics and AI. ToddlerBot enables seamless acquisition of high-quality simulation and real-world data. The plug-and-play zero-point calibration and transferable motor system identification ensure a high-fidelity digital twin, enabling zero-shot policy transfer from simulation to the real-world. A user-friendly teleoperation interface facilitates streamlined real-world data collection for learning motor skills from human demonstrations. Utilizing its data collection ability and anthropomorphic design, ToddlerBot is an ideal platform to perform whole-body loco-manipulation. Additionally, ToddlerBot’s compact size ($0.56\leavevmode\nobreak\ \mathrm{m},3.4\leavevmode\nobreak\ \mathrm{kg}$) ensures safe operation in real-world environments. Reproducibility is achieved with an entirely 3D-printed, open-source design and commercially available components, keeping the total cost under $6000\leavevmode\nobreak\ \mathrm{USD}$. Comprehensive documentation allows assembly and maintenance with basic technical expertise, as validated by a successful independent replication of the system. We demonstrate ToddlerBot’s capabilities through arm span, payload, endurance tests, loco-manipulation tasks, and a collaborative long-horizon scenario where two robots tidy a toy session together. By advancing ML-compatibility, capability, and reproducibility, ToddlerBot provides a robust platform for scalable learning and dynamic policy execution in robotics research.
Figure 1: ToddlerBot is an open-source, ML-compatible humanoid platform for efficiently collecting large-scale, high-quality training data in both simulation and the real world. We leverage massive parallel environments, an accurate digital twin for simulation, and an intuitive teleoperation device for precise upper and lower body control in real-world data collection. We demonstrate that ToddlerBot meets these requirements and successfully acquires a diverse set of loco-manipulation skills from both data sources, including walking, push-ups, pull-ups, wagon pushing, bimanual, and full-body manipulation.
I Introduction
Humanoid Size(m) Weight(kg) # Active DoFs(a) Manipulation Locomotion Sim Data Real Data Open Source Price(\$)
BD Atlas [8] 1.50 89.0 28 - - -
Berkeley [27] 0.85 16.0 12 code 10K
Booster T1 [35] 1.18 30.0 23 code 34K
BRUCE [28] 0.70 4.8 16 code 6.5K
Cassie [33] 1.15 35.0 10 code 250K
Digit [34] 1.75 65.0 16 code 250K
Duke [49] 1.00 30.0 10 everything 16K
Figure [9] 1.68 70.0 26 - - -
Fourier GR1 [10] 1.65 55.0 32 - - 110K
iCub [32] 1.04 24.0 32 code 300K
MIT [5] 1.04 24.0 18 - - -
NAO H25 [31] 0.57 5.2 23 code 14K
Optimus [44] 1.73 57.0 28 - - -
Robotis OP3 [36] 0.51 3.5 20 code 11K
Unitree G1 [47] 1.32 35.0 29 code 57K
Unitree H1 [46] 1.76 47.0 19 code 70K
Ours 0.56 3.4 30 everything(b) 6K
Average Adult [14] 1.73 70.9 32(c) - - - -
  • (a)
    The active degrees of freedom actuated by motors, excluding end effectors such as parallel grippers or dexterous hands.
  • (b)
    Everything includes the digital twin, learning algorithms, hardware design, assembly manual, detailed documentation, and tutorials.
  • (c)
    While human body is powered by over 600 muscles in reality, the primary functional movements of the human body can be approximated using 32 revolute joints: six DoFs per leg, seven DoFs per arm, three DoFs for the waist, and three DoFs for the neck, excluding fingers and toes.
Table I: Comparison with Other Humanoid Research Platforms.
Conventional robot design prioritizes factors such as actuator strength, sensor accuracy, mechanical precision, and repeatability—key objectives for developing robust control algorithms. However, these platforms are not inherently aligned with modern robot learning paradigms driven by embodied data. A robot platform compatible with a machine learning (ML) approach must possess the innate ability to collect observation and action data seamlessly, both in simulation and in the real world, as these complementary data sources are essential for scalable policy learning. Simulation enables rapid and scalable data collection but relies heavily on accurate physics models. In contrast, real-world data tend to be more reliable but are often difficult to scale due to cost and safety concerns.
While some recent quadrupeds [23, 24] and robotic manipulators [50, 48, 41, 38, 2] have been designed with ML-compatibility in mind, we introduce ToddlerBot, a humanoid robot platform for robotics and AI research community (Figure 1), specifically developed to facilitate policy learning for both locomotion and manipulation skills. ToddlerBot is designed to maximize ML-compatibility, capability, and reproducibility, while minimizing the expertise and costs required for its construction and maintenance.
To address the need for ML-compatibility, ToddlerBot is designed not only to execute policies but also to serve as a robust data collection platform. It enables the acquisition of high-quality simulation data through a plug-and-play zero-point calibration procedure and transferable motor system identification (sysID) results. These tools ensure a high-fidelity digital twin without the need for additional tuning. We validate the quality of simulation data with both keyframe-interpolated motions (e.g., push-ups and pull-ups) and reinforcement learning (RL) policies (e.g., walking and turning), demonstrating the capability for zero-shot sim-to-real transfer. ToddlerBot can also acquire scalable real-world data. We design an intuitive teleoperation interface that allows simultaneous control of ToddlerBot’s upper and lower body to collect whole-body manipulation data and develop effective visuomotor policies. Additionally, ToddlerBot’s small size and weight ($0.56\leavevmode\nobreak\ \mathrm{m},3.4\leavevmode\nobreak\ \mathrm{kg}$) ensure safe and accessible operation in real-world environments.
Beyond ML-compatibility, ToddlerBot is designed with a focus on capability and reproducibility. A humanoid has the potential to utilize scalable human demonstrations by leveraging its anatomical similarity to the human body. As such, ToddlerBot features an anthropomorphic design with 30 active degrees of freedom (excluding end effectors), powered by carefully selected motors comparable to human muscle strengths normalized by the body size. We demonstrate ToddlerBot’s capabilities through arm span, payload, and endurance tests and various open-loop and closed-loop loco-manipulation tasks.
Reproducibility is achieved through low-cost, open-source designs and readily accessible hardware components. ToddlerBot uses commercially available motors and is completely 3D-printed, with a total cost under $6,000\leavevmode\nobreak\ \mathrm{USD}$ ($90\%$ of the cost is for motors and computers). We will release digital twin software, learning algorithms, hardware designs, and comprehensive tutorials to ensure that ToddlerBot can be built at home with basic knowledge of hardware and software, without requiring specialized equipment for manufacturing or repair. To validate reproducibility, we enlisted a CS-major student who is not involved in this project to independently build another instance of ToddlerBot with the provided assembly manual and successfully zero-shot transferred loco-manipulation policies between the two instances. Finally, we showcase a collaborative long-horizon task in which two ToddlerBot robots work together to tidy a room by organizing stuffed toys from a table and the floor.
II Related Works
In recent years, numerous humanoid robots have been developed, showcasing diverse designs and capabilities. Industrial humanoids such as Boston Dynamics Atlas [8], Booster T1 [35], Cassie [33], Digit [34], Figure [9], Fourier GR1 [10], NAO H25 [31], Tesla Optimus [44], Robotis OP3 [36], Unitree G1 [47], and H1 [46] demonstrate remarkable technological advancements. On the other hand, humanoids from research institutions, including the Berkeley Humanoid [27], BRUCE [28], Duke Humanoid [49], iCub [32], and MIT Humanoid [5], also explore the humanoid design space with different emphasis. Humanoid robots can be evaluated using various metrics; we prioritize three core attributes: ML-compatibility, capability, and reproducibility. As outlined in Table I, we identify nine metrics to compare these attributes across different humanoid platforms.
Size and weight are critical factors when designing humanoid robots. A smaller humanoid is inherently cheaper, easier to build and repair, and safer. Larger, full-size humanoids typically require a substantial engineering team for operation and maintenance, along with specialized facilities like gantry cranes for safety. In contrast, smaller humanoids can be deployed by a small team, often a single person, and operated in constrained environments with simply a laptop. But what is the trade-off with a smaller size? Reduced size can limit a robot’s ability to manipulate human-scale objects, However, miniature humanoids can still effectively perform meaningful manipulation tasks when paired with appropriately scaled objects. Furthermore, learning to manipulate smaller objects and developing dynamic whole-body control and locomotion techniques are likely transferable to larger humanoids.
We argue that the number of active DoFs is crucial for the research value of a humanoid platform since more DoFs enable more human-like motion. While the human musculoskeletal system employs over 600 muscles working in complex synergies, the primary functional movements of the human body can be approximated using 32 revolute joints in a robotic system: six DoFs per leg, seven DoFs per arm, three DoFs for the waist, and three DoFs for the neck, excluding fingers and toes. Therefore, humanoid designs aim to achieve a DoF count as close to 32 as possible. The passive DoFs are less important as they do not contribute to the action space. We believe that the public perceived limited performance of miniature humanoids is primarily due to fewer DoFs. This limitation often arises from space constraints that restrict the incorporation of many DoFs, a challenge we have successfully addressed in ToddlerBot.
To qualitatively assess the capability, we evaluate the humanoid’s ability to perform both manipulation and locomotion tasks. While each type of motion is important, the combination of both is particularly compelling, as it unlocks opportunities for whole-body control research [19, 20, 12, 22, 29]. Furthermore, certain motions, such as push-ups, pull-ups, and cartwheels, go beyond traditional categories of manipulation and locomotion, treating them as prerequisites and requiring coordinated use of both arms and legs.
Recent advances show that large-scale simulation-based data collection is highly effective for locomotion [39, 42, 26], while real-world data collection is more promising for manipulation [6, 25, 30]. Therefore, an ideal humanoid research platform should facilitate data collection in both simulation and real-world settings, which we defined as ML-compatibility.
Moreover, being open-source and low-cost is essential for others to reproduce. Without these qualities, research in this field would remain restricted to those with specialized expertise and significant resources. While making no compromise in functionality, ToddlerBot stands out as completely open-source and the most affordable among recent humanoid platforms, making it accessible to a wider range of researchers.
III System Design
Figure 2: Humanoid Metrics. Two key criteria of humanoid capability are the number of active DoFs and power factor $\tilde{p}$ (Equation 1). The total power factor is the sum of the upper and lower body power factor: $\tilde{p}_{\text{total}}=\tilde{p}_{\text{upper}}+\tilde{p}_{\text{lower}}$. ToddlerBot is the closest to human compared with other humanoids, implying potentially comparable loco-manipulation capabilities.
Figure 3: Mechatronic Design. We present three main sections: arm, leg, and head plus torso. We highlight ToddlerBot’s 30 active DoFs with orange markers: 7 DoFs per arm, 6 DoFs per leg, a 2-DoF neck, and a 2-DoF waist. Green markers indicate two end-effector designs—a compliant palm and a parallel-jaw gripper. Purple markers denote the sensor and electronics layout with exploded views, featuring two fisheye cameras, a speaker, two microphones, an IMU, and a Jetson Orin NX computer.
When designing humanoids, numerous factors must be considered, including payload, endurance, repeatability, degrees of freedom, cost, locomotion, and manipulation capabilities. Given this vast design space, ToddlerBot’s key design principles prioritize reproducibility, capability, and ML-compatibility. Unlike most prior works, we treat reproducibility as a hard constraint to narrow the design space, as our humanoid platform holds no value if it cannot be reproduced by others. With this assumption, we optimize the hardware for capability and enhance ML-compatibility through the careful design of the digital twin and teleoperation device.
III-A Reproducibility - A Hard Constraint
We define reproducibility as the ability of a single person to replicate the robot system at home without specialized equipment. Limited access to manpower and facilities presents several challenges, including fabrication, space, and cost.
Fabrication. 3D printing has become a popular method for reproducing open-source hardware systems due to its accessibility and fast turnaround time [24, 41, 48]. Despite these merits, a key challenge is to ensure that printed parts are strong enough for a humanoid robot. We find that PLA suits most parts, while high-strength components can be reinforced with a stronger printing profile and carbon fiber-blended filaments. To ensure ease of assembly and maintenance, we detail our design considerations in Section VIII-A of the supplementary material. For non-3D-printable components, such as motors, we limit ourselves to readily available off-the-shelf items.
Space. In a home environment without equipment to ensure safety, it is essential to minimize the robot’s size. However, the robot’s size is constrained by the onboard computer required for policy execution. We use the NVIDIA Jetson Orin NX for AI model execution, and to fit both the computer and battery within the torso while maintaining human-like proportions, the robot’s height is capped at approximately $0.6\leavevmode\nobreak\ \mathrm{m}$.
Cost. With 3D printing, the total BOM cost of ToddlerBot is $6000\leavevmode\nobreak\ \mathrm{USD}$, with 90% spent on the computers and motors.
III-B Capability - Design Objective
To quantitatively assess a humanoid robot’s capability, we propose two key metrics: power factor and the number of active DoFs. The power factor is defined as:
$\tilde{\mathit{p}}=\frac{\sum_{i=0}^{\mathit{N}}|{\tau}^{\max}_{i}|}{{h}\cdot{mg}},$ (1)
where $\mathit{N}$ denotes the number of active DoFs, $|{\tau}_{i}^{\max}|$ represents the peak torque of the $i$-th joint motor, $h$ denotes the humanoid’s height, and $mg$ represents the weight. We define $\tilde{\mathit{p}}$ as the power factor, representing the total torque (and thus mechanical power) a robot can generate relative to its weight and height scale. A detailed discussion of this metric is provided in Section VIII-B of the supplementary material.
Intuitively, having a higher $\tilde{\mathit{p}}$ means that a humanoid can perform energetic, dynamic motions more easily. We argue that $\tilde{\mathit{p}}$ should at least exceed the human threshold $\tilde{\mathit{p}}_{\text{human}}$ to achieve human-like motion, given the inherent gap between robot and human policies, assuming humans operate as an oracle policy that is energy efficient. However, raising $\tilde{\mathit{p}}$ far beyond $\tilde{\mathit{p}}_{\text{human}}$ can have adverse effects: it may lead to unnatural motion, excessive reliance on motor power, fewer DoFs to accommodate larger motors, reduced battery life, and increased safety concerns. Thus, pushing $\tilde{\mathit{p}}$ past diminishing returns involves a practical trade-off. As shown in Figure 2, ToddlerBot has a $\tilde{\mathit{p}}$ score closest to humans.
Mechatronic Design. As shown in Figure 3, ToddlerBot replicates the structure of an adult human body to maximize the number of active DoFs, enabling human-like motion and versatile loco-manipulation tasks. Each arm features seven DoFs with spur gears for axis-aligned transmission, ensuring high functionality and reachability. Each leg has six DoFs: three at the hip for a wide range of motion and optimized walking strides, a parallel linkage at the knee to reduce inertia, and two at the ankle for stable locomotion. The neck, with two DoFs, incorporates parallel linkages at the pitch joint for a compact design, allowing expressive motion and full head mobility. The waist, also with two DoFs, uses coupled bevel gears to balance the space budget and effectively transmit power from two motors for yaw and roll actuation, which enables whole-body control. Additionally, ToddlerBot offers two end-effector designs: a parallel jaw gripper for grasping and a compliant palm for tasks requiring a palm-like posture. The end-effector designs can be switched quickly within two minutes by removing a few screws.
For each active DoF, we maximize the range of motion by optimizing geometries to prevent self-collisions as shown in Section VIII-C of the supplementary material. To address additional concerns such as space constraints, axis alignment, and inertia reduction, we integrate three primary transmission mechanisms in the mechanical design: spur gears, coupled bevel gears, and parallel linkages, as discussed in Section VIII-D of the supplementary material.
Motor Selection. We choose Dynamixel motors because of their robustness, reliability, and accessibility. Different types of Dynamixel motors were selected for various joints based on space constraints, torque requirements, and cost considerations. In terms of communication speed, Dynamixel motors communicate via a $5\mathrm{V}$ TTL protocol running at 2M baudrate, providing full-state feedback for all 30 motors at $50\leavevmode\nobreak\ \mathrm{Hz}$ using an off-the-shelf communication board. More detailed considerations are discussed in Section VIII-E of the supplementary material.
Sensors, Compute, and Power. To enhance ToddlerBot’s capabilities, we integrate a comprehensive set of sensors and computational components. Two fisheye cameras are included to expand the field of view, improving situational awareness. An Inertial Measurement Unit (IMU) is placed in the chest to provide state feedback for precise control. A speaker and microphones facilitate seamless communication with humans and other ToddlerBot instances. We provide a conversation example between two ToddlerBot instances in the supplementary video. The onboard computation is powered by a Jetson Orin NX 16GB, enabling real-time inference of machine learning models. Power management is handled by a custom-designed power distribution board with dual $12\mathrm{V}$ converters. This system efficiently regulates the $12\mathrm{V}$ rail and distributes power to various components, meeting the high energy demands while fitting within the limited space of the torso. The power budget is discussed in Section VIII-F of the supplementary material.
Figure 4: Zero-point Calibration. We 3D-print devices for the plug-and-play zero-point calibration procedure: orange for the arm, yellow for the neck, red for the hip, and beige for the ankle. Arrows indicate the insertion direction, and the zero-point is fixed once the devices click into place.
III-C ML-Compatibility - Design Objective
Digital Twin. A high-fidelity digital twin is essential for high-quality simulation data collection and thus zero-shot sim-to-real transfer. We divide the digital twin development into two key components: zero-point calibration for correct kinematics and motor system identification for accurate dynamics.
Since Dynamixel motors lack an inherent zero point, a reliable method is needed to recalibrate after assembly, which is frequent during repairs or design iterations. As shown in Figure 4, we design calibration devices in CAD that quickly align the robot to its zero point, defined as standing with both arms besides the body. The process takes less than a minute.
With kinematics correctly calibrated, the next challenge is accurately identifying dynamics parameters. Firstly, while it may seem trivial, it is essential to carefully weigh all components to ensure an accurate weight distribution for the robot. Inspired by Haarnoja et al. [17], we collect sysID data by commanding the motors to track a chirp signal and use the resulting position tracking data to fit an actuation model as described in Grandia et al. [13]. The actuation model is detailed in Section VIII-G of the supplementary material. We assume that motors of the same model have nearly identical dynamics parameters due to manufacturing consistency. This assumption is empirically validated by performing sysID only five times—once for each Dynamixel motor model—and successfully transferring policies to a second ToddlerBot instance without additional sysID.
Teleoperation Device. Previous works [1, 48] demonstrate that teleoperation is one of the most effective ways to gather real-world data. Inspired by them, we develop a second upper body of ToddlerBot as the leader arms. Two force-sensitive resistors (FSRs) are embedded in the gripping area of the end effectors to detect compression force from the operator, allowing gripper movement based on force input.
We use a handheld gaming computer (either Steam Deck or ROG Ally X) to control the other body parts. The joysticks send velocity commands to walk, turn, and squat. Buttons trigger either programmed or trained policies, where holding a button executes the policy continuously, and releasing it stops the action. Buttons also provide direct control over neck and waist movements. The detailed mapping is described in Section VIII-H of the supplementary material.
IV System Control
IV-A Keyframe Animation
Keyframe animation is a cornerstone of character animation, but it provides only kinematic data, with no guarantee of dynamic feasibility [21]. To address this, we developed software integrating MuJoCo [45] with a GUI, enabling real-time tuning and validation of keyframes and motion trajectories generated through linear interpolation with user-defined timings. With our high-fidelity digital twin, we efficiently generate open-loop trajectories such as cuddling, push-ups, and pull-ups that can be executed zero-shot in the real world.
IV-B Reinforcement Learning
For walking and turning, we train a reinforcement learning (RL) policy, $\pi(\bm{\mathrm{a}}_{t}|\bm{\mathrm{s}}_{t})$, which outputs $\bm{\mathrm{a}}_{t}$ as joint position setpoints for proportional-derivative (PD) controllers, based on the observable state $\bm{\mathrm{s}}_{t}$:
$\bm{\mathrm{s}}_{t}=\left(\bm{\phi}_{t},\bm{\mathrm{c}}_{t},\Delta\bm{\mathrm{q}}_{t},\bm{\dot{\mathrm{q}}}_{t},\bm{\mathrm{a}}_{t-1},\bm{\theta}_{t},\bm{\omega}_{t}\right),$ (2)
where $\bm{\phi}_{t}$ is a phase signal, $\bm{c}_{t}$ represents velocity commands, $\Delta\bm{q}_{t}$ denotes the position offset relative to the neutral pose $\bm{q}_{0}$, $\bm{a}_{t-1}$ is the action from the previous time step, $\bm{\theta}_{t}$ represents the torso orientation, and $\bm{\omega}_{t}$ is the torso’s angular velocity. All inputs are normalized to ensure stable learning.
During PPO policy training [40], the environment generates the next state, $\bm{s}_{t+1}$, updates the phase signal, and returns a scalar reward $\mathrm{r}_{t}=\mathrm{r}(\bm{s}_{t},\bm{a}_{t},\bm{s}_{t+1},\bm{\phi}_{t},\bm{c}_{t})$. Following standard practice [13], the reward is decomposed as:
$\mathrm{r}_{t}=\mathrm{r}_{t}^{\text{imitation}}+\mathrm{r}_{t}^{\text{regularization}}+\mathrm{r}_{t}^{\text{survival}}.$ (3)
Among these components, $\mathrm{r}_{t}^{\text{imitation}}$ encourages accurate imitation of the reference walking motion, which is generated using a closed-form ZMP (Zero Moment Point) solution [43]. $\mathrm{r}_{t}^{\text{regularization}}$ incorporates heuristics of ideal walking motion, penalizes joint torques, and promotes smooth actions to minimize unnecessary movements. A survival reward $\mathrm{r}_{t}^{\text{survival}}$ prevents early episode termination during training. Additional details are provided in Section VIII-I of the supplementary material.
Figure 5: Arm Span and Payload. We show that with a torso dimension of $13\times 9\times 12\leavevmode\nobreak\ \mathrm{cm}^{3}$, ToddlerBot can grasp objects up to $27\times 24\times 31\leavevmode\nobreak\ \mathrm{cm}^{3}$, about 14 times the torso size. Additionally, ToddlerBot can lift weights up to $1484\leavevmode\nobreak\ \mathrm{g}$, which is 40% of its body weight ($3484\leavevmode\nobreak\ \mathrm{g}$).
IV-C Imitation Learning
Real-world data collection involves a human operator guiding the leader’s arms to teleoperate the follower’s arms while using a joystick and buttons on a remote controller to control body movements.
During data collection, when the upper body tracks the position commands from the leader arms, ToddlerBot’s lower body employs a two-layer PD controller to maintain balance. The first layer is a Center of Mass (CoM) PD controller, which keeps the CoM close to the center of the support polygon. The first layer addresses CoM shifts caused by arm movements, while the second layer manages lifting heavy objects. The second layer is a torso pitch PD controller, which uses IMU readings to ensure the torso remains upright.
Figure 6: Trajectory Tracking. We present ten consecutive real-world rollouts of an RL walking policy tracking a square trajectory with a predefined velocity profile. Both raw and smoothed linear and angular velocity tracking are displayed, with real-world results averaged across trials.
Table II: Tracking Errors in Simulation versus the Real World
Tracking Errors Simulation Real-World
Position [m] 0.082 0.133 $\pm$ 0.018
Linear Velocity [m/s] 0.016 0.032 $\pm$ 0.002
Angular Velocity [rad/s] 0.056 0.113 $\pm$ 0.010
With this setup, we can collect 60 trajectories in just 20 minutes for both bimanual and full-body manipulation tasks. The motor positions of the leader’s arms are recorded as the actions, while the motor positions of the follower robot, along with the RGB images captured from its camera, are recorded as observations. This data is subsequently used to train a diffusion policy [3], with further details provided in Section VIII-J of the supplementary material.
V Experiments
We conducted a series of experiments to show that ToddlerBot is capable, ML-compatible, and reproducible.
Figure 7: Experiment Results. We present a series of photos showcasing four different tasks: push-up, pull-up, bimanual, and full-body manipulation, demonstrating ToddlerBot’s capability in challenging locomotion and manipulation tasks.
V-A Capability: Arm Span, Payload, and Endurance
To evaluate ToddlerBot’s arm span and payload capacity, we teleoperate it to hug a test object using the compliant palm gripper while maintaining balance. In the arm span test, we show that with a torso dimension of $13\times 9\times 12\leavevmode\nobreak\ \mathrm{cm}^{3}$, ToddlerBot can grasp objects up to $27\times 24\times 31\leavevmode\nobreak\ \mathrm{cm}^{3}$, approximately 14 times its torso volume.
The payload test assesses both the upper body’s lifting capacity and the lower body’s ability to maintain balance. ToddlerBot successfully lifts up to $1484\leavevmode\nobreak\ \mathrm{g}$, 40% of its total weight ($3484\leavevmode\nobreak\ \mathrm{g}$). To eliminate friction effects, we use a 3D-printed cup for the compliant palm gripper to lock in securely. To determine the true limit, we incrementally add screws to the cup until ToddlerBot falls over. Results are shown in Figure 5.
In the endurance test, ToddlerBot starts with a fully charged battery, running the walking RL policy while stepping in place. ToddlerBot achieved the longest streak of 19 minutes without falling. Over time, increased motor temperatures gradually pushed it outside the policy’s training distribution, leading to more frequent falls. ToddlerBot withstands up to 7 falls before breaking, but even then, repairs are quick—requiring only 21 minutes of 3D printing and 14 minutes of assembly, including removing the damaged part, installing the replacement, performing zero-point calibration, and rerunning the script.
V-B Capability: Push-ups and Pull-ups
To demonstrate ToddlerBot’s ability to execute expressive and dynamic motions, we program push-ups and pull-ups in our keyframe software and perform a zero-shot sim-to-real transfer. For pull-ups, we use an AprilTag to help ToddlerBot accurately locate the horizontal bar. Both tasks require strong limbs, balanced upper-lower body strength, and precise coordination, particularly when ToddlerBot transitions from a planking pose to standing after push-ups and when it releases the horizontal bar and lands after pull-ups. These open-loop transfers require only a single motion trajectory designed in simulation, highlighting the fidelity of our digital twin. Figure 7 showcases both tasks through a series of images.
Figure 8: Skill Chaining. We demonstrate that ToddlerBot can seamlessly transition from a DP-based grasping skill to an RL-trained walking policy, enabling it to grasp the wagon handle and push it forward while maintaining its grip.
Figure 9: Two-instance Collaboration. In this task, two instances of ToddlerBot, Arya and Toddy, collaborate to clean up a toy session. (1) The scene begins with a pink octopus on the table, a purple octopus on the ground, a wagon, and Arya standing next to the table. (2) Arya picks up the pink octopus from the table and places it in the wagon. (3) Arya walks to the wagon handle. (4) Arya runs a grasping policy to grasp the wagon handle, while Toddy walks over. (5) Arya pushes the wagon toward the purple octopus and waits for Toddy. (6) Toddy reaches the pickup position. (7) Toddy kneels and picks up the purple octopus. (8) Finally, Arya and Toddy walk side by side, leaving the scene together.
V-C ML-Compatibility: Omnidirectional Walking
To demonstrate ToddlerBot’s locomotion capabilities, we train RL walking policies to follow a square trajectory with a predefined velocity profile. Figure 6 and Table II present the results, with real-world tracking data collected via motion capture. Due to RL policy limitations, both simulation and real-world tracking deviate from the command, primarily because the learned walking policy struggles with in-place rotation, causing translation offsets. However, the sim-to-real gap is notably smaller than the tracking gap, supporting a successful zero-shot transfer. Additionally, we report a position tracking error variance of $0.018\leavevmode\nobreak\ \mathrm{m}$, linear velocity tracking error of $0.002\leavevmode\nobreak\ \mathrm{m/s}$, and angular velocity tracking error of $0.01\leavevmode\nobreak\ \mathrm{rad/s}$, which demonstrates good repeatability.
V-D ML-Compatibility: Vision-based Manipulation
We show ToddlerBot’s ability to perform bimanual manipulation and full-body manipulation by transferring octopus toys from a table and the ground to a wagon. Both tasks are trained with RGB-based diffusion policy [3] with 60 demonstrations. Across 20 test trials, we achieve a 90% success rate for bimanual manipulation and 75% for full-body manipulation. We leverage a combination of open-loop motions and closed-loop policies to enhance data collection efficiency. In the bimanual task, the torso rotating and releasing motions are open-loop, while in full-body manipulation, kneeling down is open-loop. The same motions are also used during policy evaluation. ToddlerBot’s onboard computing runs a 300M-parameter diffusion policy with about $100\leavevmode\nobreak\ \mathrm{ms}$ latency, enabling real-time operation. Results are shown in Figure 7.
V-E ML-Compatibility: Skill Chaining
To test the system’s ability to combine locomotion and manipulation skills, we test the wagon-pushing task (Figure 8), which requires the robot to first perform vision-based grasping and then walk forward. To do so, ToddlerBot first executes a diffusion policy to grasp the handle, while maintaining that pose, switched to the RL policy to push the wagon forward. To enable ToddlerBot to walk while maintaining its grip, we sample the end pose of the upper body from the grasping policy’s training data (60 demonstrations) during RL training.
V-F Reproducibility: Hardware and Policies
To demonstrate hardware reproducibility, we recruit a CS-major student with no prior hardware experience to assemble a second ToddlerBot using our manuals as shown in Figure 15 of the supplementary material and our open-source assembly videos. The student independently completes the assembly in three days, including the time required for 3D printing.
For policy reproducibility, we run the manipulation policy trained on data collected with one instance on the other instance, achieving the same success rate of 90% across 20 trials. We also successfully transfer the RL walking policy between both robots. To further showcase the equivalent performance of both ToddlerBot instances, we have them collaborate on a long-horizon room tidying task, as shown in Figure 9.
VI Limitations and Future Work
ToddlerBot’s performance in more agile tasks is constrained by the off-the-shelf motors’ max speed, max torque, and communication speed. Rather than achieving superhuman capabilities, ToddlerBot aligns more closely with average human performance in loco-manipulation tasks. Additionally, its scale limits interaction with human-sized objects, though this does not hinder research if appropriately sized objects are used.
To overcome these limitations, we are developing customized communication boards and improving motor system identification to maximize performance. We also aim to improve sensing capabilities, including stereo vision for depth perception, additional IMUs for improved state estimation, and tactile sensors for richer feedback.
VII Conclusions
In conclusion, we demonstrate that ToddlerBot is ML-compatible, capable, and reproducible through a series of tests and loco-manipulation tasks. While humanoid research is often associated with locomotion, ToddlerBot extends beyond this to support full-body manipulation, character animation, human-robot interaction, and various ML applications, making it a versatile research platform. With ToddlerBot fully open-source, we hope to empower researchers to explore new directions in humanoid research and encourage open collaboration in the community.
Acknowledgments
The authors would like to express their gratitude to Kaizhe Hu for assembling the second instance of ToddlerBot and assisting with keyframe animation and demo recording. We also extend our thanks to Huy Ha, Yen-Jen Wang, Pei Xu, and Yifan Hou for their insightful discussions on locomotion, and to Sirui Chen, Chen Wang, and Yunfan Jiang for valuable input on manipulation policy deployment. We are grateful to Albert Wu for his guidance on mathematical formulation and notation. Additionally, we thank João Pedro Araújo for his assistance with the motion capture system. Finally, we appreciate the helpful discussions from all members of TML and REALab. This work was supported by National Science Foundation NSF-FRR-2153854, NSF-2143601, NSF-2037101, Sloan Fellowship, Stanford Institute for Human-Centered Artificial Intelligence, and Stanford Wu Tsai Human Performance Alliance.

### v3

@@
 Abstract
-Learning-based robotics research driven by data demands a new approach to robot hardware design—one that serves as both a platform for policy execution and a tool for embodied data collection to train policies.
-We introduce ToddlerBot, a low-cost, open-source humanoid robot platform designed for scalable policy learning and research in robotics and AI.
+Learning-based robotics research driven by data demands a new approach to robot hardware design—one that serves as both a platform for policy execution and a tool for embodied data collection.
+We introduce ToddlerBot, a low-cost, open-source humanoid robot platform designed for robotics and AI research.
 ToddlerBot enables seamless acquisition of high-quality simulation and real-world data.
-The plug-and-play zero-point calibration and transferable motor system identification ensure a high-fidelity digital twin, enabling zero-shot policy transfer from simulation to the real-world.
-A user-friendly teleoperation interface facilitates streamlined real-world data collection for learning motor skills from human demonstrations.
-Utilizing its data collection ability and anthropomorphic design, ToddlerBot is an ideal platform to perform whole-body loco-manipulation.
+The plug-and-play zero-point calibration and transferable motor system identification ensure a high-fidelity digital twin and zero-shot sim-to-real policy transfer.
+A user-friendly teleoperation interface streamlines real-world data collection from human demonstrations.
+With its data collection ability and anthropomorphic design, ToddlerBot is ideal for whole-body loco-manipulation research.
 Additionally, ToddlerBot’s compact size ($0.56\leavevmode\nobreak\ \mathrm{m},3.4\leavevmode\nobreak\ \mathrm{kg}$) ensures safe operation in real-world environments.
-Reproducibility is achieved with an entirely 3D-printed, open-source design and commercially available components, keeping the total cost under $6000\leavevmode\nobreak\ \mathrm{USD}$.
-Comprehensive documentation allows assembly and maintenance with basic technical expertise, as validated by a successful independent replication of the system.
+Reproducibility is achieved with entirely 3D-printed, open-source design and off-the-shelf components, keeping the total cost under $6,000\leavevmode\nobreak\ \mathrm{USD}$.
+This allows assembly and maintenance with basic technical expertise, as validated by successful independent replications of the system.
 We demonstrate ToddlerBot’s capabilities through arm span, payload, endurance tests, loco-manipulation tasks, and a collaborative long-horizon scenario where two robots tidy a toy session together.
-By advancing ML-compatibility, capability, and reproducibility, ToddlerBot provides a robust platform for scalable learning and dynamic policy execution in robotics research.
-Figure 1: ToddlerBot is an open-source, ML-compatible humanoid platform for efficiently collecting large-scale, high-quality training data in both simulation and the real world.
-We leverage massive parallel environments, an accurate digital twin for simulation, and an intuitive teleoperation device for precise upper and lower body control in real-world data collection.
-We demonstrate that ToddlerBot meets these requirements and successfully acquires a diverse set of loco-manipulation skills from both data sources, including walking, push-ups, pull-ups, wagon pushing, bimanual, and full-body manipulation.
-I Introduction
-Humanoid Size(m) Weight(kg) # Active DoFs(a) Manipulation Locomotion Sim Data Real Data Open Source Price(\$)
-BD Atlas [8] 1.50 89.0 28 - - -
-Berkeley [27] 0.85 16.0 12 code 10K
-Booster T1 [35] 1.18 30.0 23 code 34K
-BRUCE [28] 0.70 4.8 16 code 6.5K
-Cassie [33] 1.15 35.0 10 code 250K
-Digit [34] 1.75 65.0 16 code 250K
-Duke [49] 1.00 30.0 10 everything 16K
-Figure [9] 1.68 70.0 26 - - -
-Fourier GR1 [10] 1.65 55.0 32 - - 110K
-iCub [32] 1.04 24.0 32 code 300K
-MIT [5] 1.04 24.0 18 - - -
-NAO H25 [31] 0.57 5.2 23 code 14K
-Optimus [44] 1.73 57.0 28 - - -
-Robotis OP3 [36] 0.51 3.5 20 code 11K
-Unitree G1 [47] 1.32 35.0 29 code 57K
-Unitree H1 [46] 1.76 47.0 19 code 70K
-Ours 0.56 3.4 30 everything(b) 6K
-Average Adult [14] 1.73 70.9 32(c) - - - -
-  • (a)
-    The active degrees of freedom actuated by motors, excluding end effectors such as parallel grippers or dexterous hands.
-  • (b)
-    Everything includes the digital twin, learning algorithms, hardware design, assembly manual, detailed documentation, and tutorials.
-  • (c)
-    While human body is powered by over 600 muscles in reality, the primary functional movements of the human body can be approximated using 32 revolute joints: six DoFs per leg, seven DoFs per arm, three DoFs for the waist, and three DoFs for the neck, excluding fingers and toes.
-Table I: Comparison with Other Humanoid Research Platforms.
-Conventional robot design prioritizes factors such as actuator strength, sensor accuracy, mechanical precision, and repeatability—key objectives for developing robust control algorithms.
-However, these platforms are not inherently aligned with modern robot learning paradigms driven by embodied data.
-A robot platform compatible with a machine learning (ML) approach must possess the innate ability to collect observation and action data seamlessly, both in simulation and in the real world, as these complementary data sources are essential for scalable policy learning.
-Simulation enables rapid and scalable data collection but relies heavily on accurate physics models.
-In contrast, real-world data tend to be more reliable but are often difficult to scale due to cost and safety concerns.
-While some recent quadrupeds [23, 24] and robotic manipulators [50, 48, 41, 38, 2] have been designed with ML-compatibility in mind, we introduce ToddlerBot, a humanoid robot platform for robotics and AI research community (Figure 1), specifically developed to facilitate policy learning for both locomotion and manipulation skills.
-ToddlerBot is designed to maximize ML-compatibility, capability, and reproducibility, while minimizing the expertise and costs required for its construction and maintenance.
-To address the need for ML-compatibility, ToddlerBot is designed not only to execute policies but also to serve as a robust data collection platform.
+By advancing ML-compatibility, capability, and reproducibility, ToddlerBot provides a robust and scalable platform for policy learning and execution in robotics research.
+Keywords: Humanoid, Mechanisms & Design, Robot Modeling & Simulation
+Figure 1: ToddlerBot is an open-source humanoid platform for large-scale, high-quality data collection in both simulation and the real world.
+It combines massive parallel simulation and an accurate digital twin for simulation, and an intuitive teleoperation device for whole-body control in the real world.
+ToddlerBot enables diverse loco-manipulation skills, including walking, push-ups, pull-ups, wagon pushing, bimanual, and full-body manipulation, learned from both data sources.
+1 Introduction
+Conventional robot design prioritizes actuator strength, sensor accuracy, and mechanical precision.
+While these characteristics are essential for deploying robots in real-world environments, they fall short in addressing the needs of robotics research—where affordability, rapid repairability, and full-stack ownership with no black boxes are crucial for faster innovation, easier debugging, and total system understanding.
+Additionally, traditional robot platforms are often not aligned with modern machine learning (ML) paradigms driven by embodied data.
+A robot platform compatible with ML must possess the innate ability to collect observation and action data, both in simulation and the real world, as these complementary data sources are essential for scalable policy learning.
+In response to these challenges, we introduce ToddlerBot, an open-source humanoid robot platform for the robotics and AI research community (Figure 1), specifically developed to facilitate policy learning for locomotion and manipulation skills.
+ToddlerBot is designed to resemble a miniature human to minimize construction and maintenance costs and specialized expertise, lowering the barrier for researchers seeking to build self-sufficient, fully transparent humanoids free from critical reliance on third-party components.
+Although ToddlerBot lacks the payload capacity of full-sized humanoids, it offers more degrees of freedom in both the upper and lower body than any existing miniature humanoids—sufficient for advancing research in manipulation and locomotion.
+Inspired by recent quadrupeds [1, 2] and robotic manipulators [3, 4, 5, 6, 7], ToddlerBot is designed with ML-compatibility in mind.
+Researchers can use ToddlerBot to not only execute policies but also to serve as a robust data collection platform.
 It enables the acquisition of high-quality simulation data through a plug-and-play zero-point calibration procedure and transferable motor system identification (sysID) results.
-These tools ensure a high-fidelity digital twin without the need for additional tuning.
-We validate the quality of simulation data with both keyframe-interpolated motions (e.g., push-ups and pull-ups) and reinforcement learning (RL) policies (e.g., walking and turning), demonstrating the capability for zero-shot sim-to-real transfer.
+These tools ensure a high-fidelity digital twin.
+We validate the quality of simulation data with keyframe-interpolated motions (e.g., push-ups and pull-ups) and reinforcement learning (RL) policies (e.g., walking), demonstrating the capability for zero-shot sim-to-real transfer.
 ToddlerBot can also acquire scalable real-world data.
@@
 Beyond ML-compatibility, ToddlerBot is designed with a focus on capability and reproducibility.
-A humanoid has the potential to utilize scalable human demonstrations by leveraging its anatomical similarity to the human body.
-As such, ToddlerBot features an anthropomorphic design with 30 active degrees of freedom (excluding end effectors), powered by carefully selected motors comparable to human muscle strengths normalized by the body size.
-We demonstrate ToddlerBot’s capabilities through arm span, payload, and endurance tests and various open-loop and closed-loop loco-manipulation tasks.
+A humanoid has the potential to utilize large scale human demonstrations by leveraging its anatomical similarity to the human body.
+As such, ToddlerBot features an anthropomorphic design with 30 active degrees of freedom (excluding end effectors).
+We demonstrate ToddlerBot’s capabilities through arm span, payload, endurance tests, and various open-loop and closed-loop loco-manipulation tasks.
 Reproducibility is achieved through low-cost, open-source designs and readily accessible hardware components.
 ToddlerBot uses commercially available motors and is completely 3D-printed, with a total cost under $6,000\leavevmode\nobreak\ \mathrm{USD}$ ($90\%$ of the cost is for motors and computers).
-We will release digital twin software, learning algorithms, hardware designs, and comprehensive tutorials to ensure that ToddlerBot can be built at home with basic knowledge of hardware and software, without requiring specialized equipment for manufacturing or repair.
-To validate reproducibility, we enlisted a CS-major student who is not involved in this project to independently build another instance of ToddlerBot with the provided assembly manual and successfully zero-shot transferred loco-manipulation policies between the two instances.
-Finally, we showcase a collaborative long-horizon task in which two ToddlerBot robots work together to tidy a room by organizing stuffed toys from a table and the floor.
-II Related Works
-In recent years, numerous humanoid robots have been developed, showcasing diverse designs and capabilities.
-Industrial humanoids such as Boston Dynamics Atlas [8], Booster T1 [35], Cassie [33], Digit [34], Figure [9], Fourier GR1 [10], NAO H25 [31], Tesla Optimus [44], Robotis OP3 [36], Unitree G1 [47], and H1 [46] demonstrate remarkable technological advancements.
-On the other hand, humanoids from research institutions, including the Berkeley Humanoid [27], BRUCE [28], Duke Humanoid [49], iCub [32], and MIT Humanoid [5], also explore the humanoid design space with different emphasis.
-Humanoid robots can be evaluated using various metrics; we prioritize three core attributes: ML-compatibility, capability, and reproducibility.
-As outlined in Table I, we identify nine metrics to compare these attributes across different humanoid platforms.
-Size and weight are critical factors when designing humanoid robots.
-A smaller humanoid is inherently cheaper, easier to build and repair, and safer.
-Larger, full-size humanoids typically require a substantial engineering team for operation and maintenance, along with specialized facilities like gantry cranes for safety.
-In contrast, smaller humanoids can be deployed by a small team, often a single person, and operated in constrained environments with simply a laptop.
-But what is the trade-off with a smaller size?
-Reduced size can limit a robot’s ability to manipulate human-scale objects, However, miniature humanoids can still effectively perform meaningful manipulation tasks when paired with appropriately scaled objects.
-Furthermore, learning to manipulate smaller objects and developing dynamic whole-body control and locomotion techniques are likely transferable to larger humanoids.
-We argue that the number of active DoFs is crucial for the research value of a humanoid platform since more DoFs enable more human-like motion.
-While the human musculoskeletal system employs over 600 muscles working in complex synergies, the primary functional movements of the human body can be approximated using 32 revolute joints in a robotic system: six DoFs per leg, seven DoFs per arm, three DoFs for the waist, and three DoFs for the neck, excluding fingers and toes.
+We release digital twin, learning algorithms, hardware designs, and comprehensive tutorials to ensure that ToddlerBot can be built at home with basic knowledge of hardware and software, without requiring specialized equipment for manufacturing or repair.
+To validate reproducibility, we enlisted a CS-major student not involved in this project to independently build another instance of ToddlerBot.
+and successfully zero-shot transfer loco-manipulation policies between the two instances.
+We also received five reports of independent replications of ToddlerBot within a week by teams worldwide.
+We summarize our contributions as follows: (1) As a system paper, we integrate existing components to create the first miniature humanoid with 30 DoFs and a superhuman range of motion, offering a practical humanoid platform with new capabilities for the community.
+(2) We develop a comprehensive sysID pipeline to ensure a high-fidelity digital twin in simulation.
+(3) We design a whole-body teleoperation solution to collect loco-manipulation data in the real world.
+2 Related Works
+Table 1: Comparison with Other Popular Humanoid Research Platforms.
+Humanoid Size Weight Compute Active DoFs(b) Manipulation Locomotion Sim Data Real Data Open Source Price
+(m) (kg) (TFLOPS)(a) - - - - - - (\$)
+BD Atlas [8] 1.50 89.0 - 28 - - -
+Figure [9] 1.68 70.0 - 26 - - -
+Optimus [10] 1.73 57.0 - 28 - - -
+Digit [11] 1.75 65.0 - 16 code 250K
+Unitree H1 [12] 1.76 47.0 1.92 19 code 70K
+Fourier GR1 [13] 1.65 55.0 2.23 32 - - 110K
+Booster T1 [14] 1.18 30.0 3.33 23 code 34K
+Unitree G1 [15] 1.32 35.0 2.50 29 code 57K
+iCub [16] 1.04 24.0 1.93 32 code 300K
+Berkeley [17] 0.85 16.0 1.92 12 code 10K
+MIT [18] 1.04 24.0 - 18 - - -
+Duke [19] 1.00 30.0 Optional 10 design, code 16K
+Berkeley Lite [20] 0.80 16.0 0.29 22 design, code 5K
+BRUCE [21] 0.70 4.8 0.1 16 code 6.5K
+NAO H25 [22] 0.57 5.2 0.02 23 code 14K
+Robotis OP3 [23] 0.51 3.5 0.30 20 code 11K
+Zeroth [24] 0.48 3.6 0.01 16 design, code 1.4K
+Ours 0.56 3.4 2.50 30 design, code 6K
+Average Adult [25] 1.73 70.9 - 32(c) - - - -
+  • (a)
+    TFLOPS refers to Tera FLoating-point Operations Per Second with single precision (FP32).
+  • (b)
+    The active degrees of freedom actuated by motors, excluding end effectors such as parallel grippers or dexterous hands.
+  • (c)
+    While human body is powered by over 600 muscles, the primary functional movements can be approximated using 32 revolute joints: six DoFs per leg, seven DoFs per arm, three DoFs for the waist, and three DoFs for the neck, excluding fingers and toes.
+In recent years, numerous humanoid robots have been developed.
+Full size industrial humanoids such as Boston Dynamics Atlas [8], Cassie [26], Digit [11], Figure [9], Tesla Optimus [10], Unitree H1 [12] and Fourier GR1 [13].
+Half-scaled Booster T1 [14], Berkeley Humanoid Lite  [20] and Unitree G1 [15].
+Miniature-sized NAO H25 [22], Robotis OP3 [23], and K-Scale Zeroth [24].
+Together, they showcase diverse designs and capabilities.
+On the other hand, humanoids from research institutions, including the Berkeley Humanoid [17], BRUCE [21], Duke Humanoid [19], iCub [16], and MIT Humanoid [18], also explore the humanoid design space with different optimization emphasis.
+As outlined in Table 1, we identify nine metrics to compare different humanoid platforms.
+Size and weight are critical when designing humanoid robots.
+For manipulation, full-size humanoids are compatible with human-scale objects, whereas reduced-size robots are limited in their ability to manipulate them.
+However, reduced-size humanoids can still benefit manipulation research when paired with appropriately scaled objects [20].
+Furthermore, dynamic whole-body control techniques are likely transferable to larger humanoids.
+For locomotion, full-sized and small-scale humanoids offer similar research values [27].
+Full-size humanoids typically require a substantial engineering team to operate and maintain, along with specialized facilities like gantry cranes for safety.
+In contrast, a smaller humanoid is inherently cheaper, easier to build and repair, and safer [28, 27, 29].
+They can be deployed by a single person and operated in constrained environments with a laptop.
+The number of active DoFs is crucial for human-like motion.
+While the human musculoskeletal system employs over 600 muscles, primary functional movement can be approximated using 32 revolute joints [16, 30]: six DoFs per leg, seven DoFs per arm, three DoFs for the waist, and three DoFs for the neck, excluding fingers and toes.
 Therefore, humanoid designs aim to achieve a DoF count as close to 32 as possible.
-The passive DoFs are less important as they do not contribute to the action space.
-We believe that the public perceived limited performance of miniature humanoids is primarily due to fewer DoFs.
+The public perception of the limited performance of miniature humanoids is primarily due to fewer DoFs.
 This limitation often arises from space constraints that restrict the incorporation of many DoFs, a challenge we have successfully addressed in ToddlerBot.
 To qualitatively assess the capability, we evaluate the humanoid’s ability to perform both manipulation and locomotion tasks.
-While each type of motion is important, the combination of both is particularly compelling, as it unlocks opportunities for whole-body control research [19, 20, 12, 22, 29].
-Furthermore, certain motions, such as push-ups, pull-ups, and cartwheels, go beyond traditional categories of manipulation and locomotion, treating them as prerequisites and requiring coordinated use of both arms and legs.
-Recent advances show that large-scale simulation-based data collection is highly effective for locomotion [39, 42, 26], while real-world data collection is more promising for manipulation [6, 25, 30].
-Therefore, an ideal humanoid research platform should facilitate data collection in both simulation and real-world settings, which we defined as ML-compatibility.
+The combination of both is particularly compelling, as it unlocks opportunities for whole-body control research [31, 32, 33, 34, 35].
+Furthermore, certain motions, such as push-ups, pull-ups, and cartwheels, require coordinated use of both arms and legs.
+Apart from physical capabilities, ToddlerBot also has best compute in class.
+Onboard CUDA accelerator allows concurrent policy inferences to support visual and locomotion capabilities.
+Large-scale simulation data collection is effective for locomotion [36, 37, 38], while real-world data collection is more promising for manipulation [39, 40, 41].
+Therefore, to be ML-compatible, a humanoid research platform should facilitate data collection in both world.
 Moreover, being open-source and low-cost is essential for others to reproduce.
 Without these qualities, research in this field would remain restricted to those with specialized expertise and significant resources.
-While making no compromise in functionality, ToddlerBot stands out as completely open-source and the most affordable among recent humanoid platforms, making it accessible to a wider range of researchers.
-III System Design
-Figure 2: Humanoid Metrics.
-Two key criteria of humanoid capability are the number of active DoFs and power factor $\tilde{p}$ (Equation 1).
-The total power factor is the sum of the upper and lower body power factor: $\tilde{p}_{\text{total}}=\tilde{p}_{\text{upper}}+\tilde{p}_{\text{lower}}$.
-ToddlerBot is the closest to human compared with other humanoids, implying potentially comparable loco-manipulation capabilities.
+While making no compromise in functionality, ToddlerBot stands out as completely open-source and the most accessible among recent humanoid research platforms.
+3 System Design
+The design space of humanoid robots is vast; ToddlerBot’s key design principles prioritize reproducibility as a hard constraint, and capability and ML-compatibility as key design objectives.
+3.1 Reproducibility - A Hard Constraint
+Unlike most prior works, we treat reproducibility as a hard constraint, as our humanoid platform holds no value if others cannot reproduce it.
+We define reproducibility as the ability of a single person to replicate the robot system at home without specialized equipment.
+3D printing has become a popular method for reproducing open-source hardware systems due to its accessibility and fast turnaround time [2, 5, 4].
+Despite these merits, it’s challenging to ensure sufficient strength and precision for printed plastic parts.
+We discuss our justifications for 3D printing in Appendix 8.1.
+For performance-critical components like motors and bearings, we limit ourselves to off-the-shelf parts.
+The total BOM cost of ToddlerBot is $6,000\leavevmode\nobreak\ \mathrm{USD}$, with 90% spent on the computer (NVIDIA Jetson) and motors (Robotis Dynamixel).
+We detail our design considerations in Appendix 8.2.
 Figure 2: Mechatronic Design.
-We present three main sections: arm, leg, and head plus torso.
-We highlight ToddlerBot’s 30 active DoFs with orange markers: 7 DoFs per arm, 6 DoFs per leg, a 2-DoF neck, and a 2-DoF waist.
+Orange markers highlights ToddlerBot’s 30 active DoFs: 7 per arm, 6 per leg, a 2 on neck, and a 2 on waist.
 Green markers indicate two end-effector designs—a compliant palm and a parallel-jaw gripper.
-Purple markers denote the sensor and electronics layout with exploded views, featuring two fisheye cameras, a speaker, two microphones, an IMU, and a Jetson Orin NX computer.
-When designing humanoids, numerous factors must be considered, including payload, endurance, repeatability, degrees of freedom, cost, locomotion, and manipulation capabilities.
-Given this vast design space, ToddlerBot’s key design principles prioritize reproducibility, capability, and ML-compatibility.
-Unlike most prior works, we treat reproducibility as a hard constraint to narrow the design space, as our humanoid platform holds no value if it cannot be reproduced by others.
-With this assumption, we optimize the hardware for capability and enhance ML-compatibility through the careful design of the digital twin and teleoperation device.
-III-A Reproducibility - A Hard Constraint
-We define reproducibility as the ability of a single person to replicate the robot system at home without specialized equipment.
-Limited access to manpower and facilities presents several challenges, including fabrication, space, and cost.
-Fabrication.
-3D printing has become a popular method for reproducing open-source hardware systems due to its accessibility and fast turnaround time [24, 41, 48].
-Despite these merits, a key challenge is to ensure that printed parts are strong enough for a humanoid robot.
-We find that PLA suits most parts, while high-strength components can be reinforced with a stronger printing profile and carbon fiber-blended filaments.
-To ensure ease of assembly and maintenance, we detail our design considerations in Section VIII-A of the supplementary material.
-For non-3D-printable components, such as motors, we limit ourselves to readily available off-the-shelf items.
-Space.
-In a home environment without equipment to ensure safety, it is essential to minimize the robot’s size.
-However, the robot’s size is constrained by the onboard computer required for policy execution.
-We use the NVIDIA Jetson Orin NX for AI model execution, and to fit both the computer and battery within the torso while maintaining human-like proportions, the robot’s height is capped at approximately $0.6\leavevmode\nobreak\ \mathrm{m}$.
-Cost. With 3D printing, the total BOM cost of ToddlerBot is $6000\leavevmode\nobreak\ \mathrm{USD}$, with 90% spent on the computers and motors.
-III-B Capability - Design Objective
-To quantitatively assess a humanoid robot’s capability, we propose two key metrics: power factor and the number of active DoFs. The power factor is defined as:
-$\tilde{\mathit{p}}=\frac{\sum_{i=0}^{\mathit{N}}|{\tau}^{\max}_{i}|}{{h}\cdot{mg}},$ (1)
-where $\mathit{N}$ denotes the number of active DoFs, $|{\tau}_{i}^{\max}|$ represents the peak torque of the $i$-th joint motor, $h$ denotes the humanoid’s height, and $mg$ represents the weight.
-We define $\tilde{\mathit{p}}$ as the power factor, representing the total torque (and thus mechanical power) a robot can generate relative to its weight and height scale.
-A detailed discussion of this metric is provided in Section VIII-B of the supplementary material.
-Intuitively, having a higher $\tilde{\mathit{p}}$ means that a humanoid can perform energetic, dynamic motions more easily.
-We argue that $\tilde{\mathit{p}}$ should at least exceed the human threshold $\tilde{\mathit{p}}_{\text{human}}$ to achieve human-like motion, given the inherent gap between robot and human policies, assuming humans operate as an oracle policy that is energy efficient.
-However, raising $\tilde{\mathit{p}}$ far beyond $\tilde{\mathit{p}}_{\text{human}}$ can have adverse effects: it may lead to unnatural motion, excessive reliance on motor power, fewer DoFs to accommodate larger motors, reduced battery life, and increased safety concerns.
-Thus, pushing $\tilde{\mathit{p}}$ past diminishing returns involves a practical trade-off.
-As shown in Figure 2, ToddlerBot has a $\tilde{\mathit{p}}$ score closest to humans.
+Purple markers denote the electronics layout with exploded views, featuring 2 fisheye cameras, 1 speaker, 2 microphones, 1 IMU, and 1 Jetson Orin computer.
+3.2 Capability - Design Objective
 Mechatronic Design.
@@
 Additionally, ToddlerBot offers two end-effector designs: a parallel jaw gripper for grasping and a compliant palm for tasks requiring a palm-like posture.
-The end-effector designs can be switched quickly within two minutes by removing a few screws.
-For each active DoF, we maximize the range of motion by optimizing geometries to prevent self-collisions as shown in Section VIII-C of the supplementary material.
-To address additional concerns such as space constraints, axis alignment, and inertia reduction, we integrate three primary transmission mechanisms in the mechanical design: spur gears, coupled bevel gears, and parallel linkages, as discussed in Section VIII-D of the supplementary material.
-Motor Selection.
-We choose Dynamixel motors because of their robustness, reliability, and accessibility.
-Different types of Dynamixel motors were selected for various joints based on space constraints, torque requirements, and cost considerations.
-In terms of communication speed, Dynamixel motors communicate via a $5\mathrm{V}$ TTL protocol running at 2M baudrate, providing full-state feedback for all 30 motors at $50\leavevmode\nobreak\ \mathrm{Hz}$ using an off-the-shelf communication board.
-More detailed considerations are discussed in Section VIII-E of the supplementary material.
+The end-effector designs can be switched quickly within two minutes.
+For each active DoF, we maximize the range of motion by optimizing geometries to prevent self-collisions (Appendix 8.3).
+To address concerns such as space constraints, axis alignment, and inertia reduction, we integrated a combination of three transmission mechanisms in the design: spur gears, coupled bevel gears, and parallel linkages (Appendix 8.4).
 Sensors, Compute, and Power.
 To enhance ToddlerBot’s capabilities, we integrate a comprehensive set of sensors and computational components.
-Two fisheye cameras are included to expand the field of view, improving situational awareness.
-An Inertial Measurement Unit (IMU) is placed in the chest to provide state feedback for precise control.
-A speaker and microphones facilitate seamless communication with humans and other ToddlerBot instances.
+Two fisheye cameras are included to expand the field of view.
+An Inertial Measurement Unit (IMU) is placed in the chest to provide state feedback.
+A speaker and microphones facilitate communication with humans and other robots.
 We provide a conversation example between two ToddlerBot instances in the supplementary video.
-The onboard computation is powered by a Jetson Orin NX 16GB, enabling real-time inference of machine learning models.
-Power management is handled by a custom-designed power distribution board with dual $12\mathrm{V}$ converters.
-This system efficiently regulates the $12\mathrm{V}$ rail and distributes power to various components, meeting the high energy demands while fitting within the limited space of the torso.
-The power budget is discussed in Section VIII-F of the supplementary material.
-Figure 4: Zero-point Calibration.
-We 3D-print devices for the plug-and-play zero-point calibration procedure: orange for the arm, yellow for the neck, red for the hip, and beige for the ankle.
-Arrows indicate the insertion direction, and the zero-point is fixed once the devices click into place.
-III-C ML-Compatibility - Design Objective
+The onboard computation is powered by a Jetson Orin NX 16GB, enabling real-time inference of ML models.
+Based on our proposed power metric (Appendix 8.5), we choose Dynamixel motors for their reliability and accessibility (Appendix 8.6).
+Power management is handled by a custom-designed power distribution board which can be easily ordered online (Appendix 8.7).
+3.3 ML-Compatibility - Design Objective
 Digital Twin.
-A high-fidelity digital twin is essential for high-quality simulation data collection and thus zero-shot sim-to-real transfer.
+A high-fidelity digital twin is fundamental for high-quality simulation data collection and zero-shot sim-to-real transfer.
 We divide the digital twin development into two key components: zero-point calibration for correct kinematics and motor system identification for accurate dynamics.
-Since Dynamixel motors lack an inherent zero point, a reliable method is needed to recalibrate after assembly, which is frequent during repairs or design iterations.
-As shown in Figure 4, we design calibration devices in CAD that quickly align the robot to its zero point, defined as standing with both arms besides the body.
-The process takes less than a minute.
-With kinematics correctly calibrated, the next challenge is accurately identifying dynamics parameters.
-Firstly, while it may seem trivial, it is essential to carefully weigh all components to ensure an accurate weight distribution for the robot.
-Inspired by Haarnoja et al.
-[17], we collect sysID data by commanding the motors to track a chirp signal and use the resulting position tracking data to fit an actuation model as described in Grandia et al.
-[13].
-The actuation model is detailed in Section VIII-G of the supplementary material.
-We assume that motors of the same model have nearly identical dynamics parameters due to manufacturing consistency.
-This assumption is empirically validated by performing sysID only five times—once for each Dynamixel motor model—and successfully transferring policies to a second ToddlerBot instance without additional sysID.
+Since Dynamixel motors lack an absolute zero point, a reliable method is needed to calibrate after reassembly.
+We designed a set of calibration devices that can align the robot to the desired zero point within a minute, defined as standing with both arms parallel to the body (Appendix 8.8).
+For accurate sysID, we designed a motor test bed and automatic test procedures described in Appendix 8.9 to measure the range of dynamics parameters.
+Next, we collect actuator position tracking data by commanding a chirp signal [27], and use the position tracking result to fit an actuation model [42] with optimization, as detailed in Appendix 8.10.
+Testing across instances shows Dynamixel motors of the same model have nearly identical dynamics parameters.
+This is also validated by successful direct deployment of policies to a second ToddlerBot instance without additional sysID.
 Teleoperation Device.
-Previous works [1, 48] demonstrate that teleoperation is one of the most effective ways to gather real-world data.
-Inspired by them, we develop a second upper body of ToddlerBot as the leader arms.
+Inspired by previous works [4, 43], we develop a second upper body of ToddlerBot as the leader arms to collect high-quality real-world data.
 Two force-sensitive resistors (FSRs) are embedded in the gripping area of the end effectors to detect compression force from the operator, allowing gripper movement based on force input.
@@
 The joysticks send velocity commands to walk, turn, and squat.
-Buttons trigger either programmed or trained policies, where holding a button executes the policy continuously, and releasing it stops the action.
-Buttons also provide direct control over neck and waist movements.
-The detailed mapping is described in Section VIII-H of the supplementary material.
-IV System Control
-IV-A Keyframe Animation
-Keyframe animation is a cornerstone of character animation, but it provides only kinematic data, with no guarantee of dynamic feasibility [21].
-To address this, we developed software integrating MuJoCo [45] with a GUI, enabling real-time tuning and validation of keyframes and motion trajectories generated through linear interpolation with user-defined timings.
-With our high-fidelity digital twin, we efficiently generate open-loop trajectories such as cuddling, push-ups, and pull-ups that can be executed zero-shot in the real world.
-IV-B Reinforcement Learning
-For walking and turning, we train a reinforcement learning (RL) policy, $\pi(\bm{\mathrm{a}}_{t}|\bm{\mathrm{s}}_{t})$, which outputs $\bm{\mathrm{a}}_{t}$ as joint position setpoints for proportional-derivative (PD) controllers, based on the observable state $\bm{\mathrm{s}}_{t}$:
+Buttons trigger either programmed or trained policies and provide direct control over neck and waist movements.
+The detailed mapping is described in Appendix 8.11.
+4 System Control
+4.1 Keyframe Animation
+Keyframe animation is widely used in character animation, but it provides only kinematics data, lacking a guarantee of dynamic feasibility [44].
+To address this, we develop a tool integrating MuJoCo [45] with a GUI, enabling real-time tuning and validation of keyframe motion trajectories.
+Combined with high-fidelity digital twin, we can efficiently generate open-loop trajectories such as cuddling, push-ups, and pull-ups that can be executed zero-shot in the real world.
+4.2 Reinforcement Learning
+For walking, we train a reinforcement learning (RL) policy, $\pi(\bm{\mathrm{a}}_{t}|\bm{\mathrm{s}}_{t})$, which outputs $\bm{\mathrm{a}}_{t}$ as joint position setpoints for proportional-derivative (PD) controllers, based on the observable state $\bm{\mathrm{s}}_{t}$:
 $\bm{\mathrm{s}}_{t}=\left(\bm{\phi}_{t},\bm{\mathrm{c}}_{t},\Delta\bm{\mathrm{q}}_{t},\bm{\dot{\mathrm{q}}}_{t},\bm{\mathrm{a}}_{t-1},\bm{\theta}_{t},\bm{\omega}_{t}\right),$ (1)
 where $\bm{\phi}_{t}$ is a phase signal, $\bm{c}_{t}$ represents velocity commands, $\Delta\bm{q}_{t}$ denotes the position offset relative to the neutral pose $\bm{q}_{0}$, $\bm{a}_{t-1}$ is the action from the previous time step, $\bm{\theta}_{t}$ represents the torso orientation, and $\bm{\omega}_{t}$ is the torso’s angular velocity.
-All inputs are normalized to ensure stable learning.
 During PPO policy training [46], the environment generates the next state, $\bm{s}_{t+1}$, updates the phase signal, and returns a scalar reward $\mathrm{r}_{t}=\mathrm{r}(\bm{s}_{t},\bm{a}_{t},\bm{s}_{t+1},\bm{\phi}_{t},\bm{c}_{t})$.
@@
 A survival reward $\mathrm{r}_{t}^{\text{survival}}$ prevents early episode termination during training.
-Additional details are provided in Section VIII-I of the supplementary material.
-Figure 5: Arm Span and Payload.
-We show that with a torso dimension of $13\times 9\times 12\leavevmode\nobreak\ \mathrm{cm}^{3}$, ToddlerBot can grasp objects up to $27\times 24\times 31\leavevmode\nobreak\ \mathrm{cm}^{3}$, about 14 times the torso size.
-Additionally, ToddlerBot can lift weights up to $1484\leavevmode\nobreak\ \mathrm{g}$, which is 40% of its body weight ($3484\leavevmode\nobreak\ \mathrm{g}$).
-IV-C Imitation Learning
-Real-world data collection involves a human operator guiding the leader’s arms to teleoperate the follower’s arms while using a joystick and buttons on a remote controller to control body movements.
-During data collection, when the upper body tracks the position commands from the leader arms, ToddlerBot’s lower body employs a two-layer PD controller to maintain balance.
+Additional details are provided in Appendix 8.12.
+4.3 Imitation Learning
+Real-world data collection involves a human operator guiding the leader’s arms to teleoperate the follower’s arms while using a joystick and buttons on a handheld game controller to control body movements.
+During data collection, when the upper body tracks the position commands from the leader arms, ToddlerBot’s lower body employs a two-layer PD controller to actively maintain balance.
 The first layer is a Center of Mass (CoM) PD controller, which keeps the CoM close to the center of the support polygon.
-The first layer addresses CoM shifts caused by arm movements, while the second layer manages lifting heavy objects.
 The second layer is a torso pitch PD controller, which uses IMU readings to ensure the torso remains upright.
-Figure 6: Trajectory Tracking.
-We present ten consecutive real-world rollouts of an RL walking policy tracking a square trajectory with a predefined velocity profile.
-Both raw and smoothed linear and angular velocity tracking are displayed, with real-world results averaged across trials.
-Table II: Tracking Errors in Simulation versus the Real World
-Tracking Errors Simulation Real-World
-Position [m] 0.082 0.133 $\pm$ 0.018
-Linear Velocity [m/s] 0.016 0.032 $\pm$ 0.002
-Angular Velocity [rad/s] 0.056 0.113 $\pm$ 0.010
+The first layer addresses CoM shifts caused by arm movements, while the second layer compensates when lifting heavy objects.
 With this setup, we can collect 60 trajectories in just 20 minutes for both bimanual and full-body manipulation tasks.
 The motor positions of the leader’s arms are recorded as the actions, while the motor positions of the follower robot, along with the RGB images captured from its camera, are recorded as observations.
-This data is subsequently used to train a diffusion policy [3], with further details provided in Section VIII-J of the supplementary material.
-V Experiments
-We conducted a series of experiments to show that ToddlerBot is capable, ML-compatible, and reproducible.
-Figure 7: Experiment Results.
-We present a series of photos showcasing four different tasks: push-up, pull-up, bimanual, and full-body manipulation, demonstrating ToddlerBot’s capability in challenging locomotion and manipulation tasks.
-V-A Capability: Arm Span, Payload, and Endurance
-To evaluate ToddlerBot’s arm span and payload capacity, we teleoperate it to hug a test object using the compliant palm gripper while maintaining balance.
-In the arm span test, we show that with a torso dimension of $13\times 9\times 12\leavevmode\nobreak\ \mathrm{cm}^{3}$, ToddlerBot can grasp objects up to $27\times 24\times 31\leavevmode\nobreak\ \mathrm{cm}^{3}$, approximately 14 times its torso volume.
+We train a diffusion policy [48] on this dataset, with further details provided in Appendix 8.13.
+5 Experiments
+Figure 3: Arm Span, Payload, and Trajectory Tracking.
+On the left, we show that with a torso dimension of $13\times 9\times 12\leavevmode\nobreak\ \mathrm{cm}^{3}$, ToddlerBot can grasp objects up to $27\times 24\times 31\leavevmode\nobreak\ \mathrm{cm}^{3}$, about 14 times the torso size.
+Additionally, ToddlerBot can lift weights up to $1484\leavevmode\nobreak\ \mathrm{g}$, which is 40% of its body weight ($3484\leavevmode\nobreak\ \mathrm{g}$).
+On the right, we present ten consecutive real-world rollouts of an RL walking policy tracking a square trajectory with a predefined velocity profile.
+Both raw and smoothed linear and angular velocity tracking are displayed, with real-world results averaged across trials.
+Capability: Arm Span, Payload, and Endurance.
+To evaluate ToddlerBot’s arm span, we teleoperate it to hug a large object using the compliant palm gripper while maintaining balance.
+We show that, ToddlerBot can grasp objects up to $27\times 24\times 31\leavevmode\nobreak\ \mathrm{cm}^{3}$, approximately 14 times its torso volume.
 The payload test assesses both the upper body’s lifting capacity and the lower body’s ability to maintain balance.
 ToddlerBot successfully lifts up to $1484\leavevmode\nobreak\ \mathrm{g}$, 40% of its total weight ($3484\leavevmode\nobreak\ \mathrm{g}$).
-To eliminate friction effects, we use a 3D-printed cup for the compliant palm gripper to lock in securely.
-To determine the true limit, we incrementally add screws to the cup until ToddlerBot falls over.
-Results are shown in Figure 5.
+To eliminate friction effects, we use a 3D-printed cup for the gripper to lock in securely.
+We incrementally add screws to the cup until ToddlerBot falls over.
+Results are shown on the left of Figure 3.
 In the endurance test, ToddlerBot starts with a fully charged battery, running the walking RL policy while stepping in place.
@@
 Over time, increased motor temperatures gradually pushed it outside the policy’s training distribution, leading to more frequent falls.
-ToddlerBot withstands up to 7 falls before breaking, but even then, repairs are quick—requiring only 21 minutes of 3D printing and 14 minutes of assembly, including removing the damaged part, installing the replacement, performing zero-point calibration, and rerunning the script.
-V-B Capability: Push-ups and Pull-ups
-To demonstrate ToddlerBot’s ability to execute expressive and dynamic motions, we program push-ups and pull-ups in our keyframe software and perform a zero-shot sim-to-real transfer.
-For pull-ups, we use an AprilTag to help ToddlerBot accurately locate the horizontal bar.
+ToddlerBot withstands up to 7 falls before breaking, but even then, repairs are quick—requiring only 21 minutes of 3D printing and 14 minutes of assembly, including removing the damaged part, installing the replacement, and performing zero-point calibration.
+Capability: Push-ups and Pull-ups.
+To demonstrate expressive and dynamic motions, we program push-ups and pull-ups in our keyframe software and perform zero-shot sim-to-real transfer (Figure 4).
+For pull-ups, we use an AprilTag to help ToddlerBot locate the horizontal bar.
 Both tasks require strong limbs, balanced upper-lower body strength, and precise coordination, particularly when ToddlerBot transitions from a planking pose to standing after push-ups and when it releases the horizontal bar and lands after pull-ups.
 These open-loop transfers require only a single motion trajectory designed in simulation, highlighting the fidelity of our digital twin.
-Figure 7 showcases both tasks through a series of images.
-Figure 8: Skill Chaining.
-We demonstrate that ToddlerBot can seamlessly transition from a DP-based grasping skill to an RL-trained walking policy, enabling it to grasp the wagon handle and push it forward while maintaining its grip.
-Figure 9: Two-instance Collaboration.
+Figure 4: Experiment Results. We present four different tasks: push-up, pull-up, bimanual, and full-body manipulation, showing ToddlerBot’s capability in challenging loco-manipulation tasks.
+Figure 5: Long-horizon Collaboration.
 In this task, two instances of ToddlerBot, Arya and Toddy, collaborate to clean up a toy session.
-(1) The scene begins with a pink octopus on the table, a purple octopus on the ground, a wagon, and Arya standing next to the table.
+(1) The task begins with a pink octopus on the table and a purple octopus on the ground.
 (2) Arya picks up the pink octopus from the table and places it in the wagon.
 (3) Arya walks to the wagon handle.
-(4) Arya runs a grasping policy to grasp the wagon handle, while Toddy walks over.
-(5) Arya pushes the wagon toward the purple octopus and waits for Toddy.
+(4) Arya grasps the handle while Toddy walks over.
+(5) Arya pushes the wagon toward the purple octopus.
 (6) Toddy reaches the pickup position.
 (7) Toddy kneels and picks up the purple octopus.
-(8) Finally, Arya and Toddy walk side by side, leaving the scene together.
-V-C ML-Compatibility: Omnidirectional Walking
+(8) Finally, Arya and Toddy leave side by side.
+ML-Compatibility: Omnidirectional Walking.
 To demonstrate ToddlerBot’s locomotion capabilities, we train RL walking policies to follow a square trajectory with a predefined velocity profile.
-Figure 6 and Table II present the results, with real-world tracking data collected via motion capture.
+The right part of Figure 3 presents the results, with real-world tracking data collected via motion capture.
 Due to RL policy limitations, both simulation and real-world tracking deviate from the command, primarily because the learned walking policy struggles with in-place rotation, causing translation offsets.
 However, the sim-to-real gap is notably smaller than the tracking gap, supporting a successful zero-shot transfer.
-Additionally, we report a position tracking error variance of $0.018\leavevmode\nobreak\ \mathrm{m}$, linear velocity tracking error of $0.002\leavevmode\nobreak\ \mathrm{m/s}$, and angular velocity tracking error of $0.01\leavevmode\nobreak\ \mathrm{rad/s}$, which demonstrates good repeatability.
-V-D ML-Compatibility: Vision-based Manipulation
+Additionally, we report a position tracking error variance of $0.018\leavevmode\nobreak\ \mathrm{m}$, which demonstrates good repeatability (more details in Appendix 8.14).
+ML-Compatibility: Vision-based Manipulation.
 We show ToddlerBot’s ability to perform bimanual manipulation and full-body manipulation by transferring octopus toys from a table and the ground to a wagon.
@@
 In the bimanual task, the torso rotating and releasing motions are open-loop, while in full-body manipulation, kneeling down is open-loop.
-The same motions are also used during policy evaluation.
-ToddlerBot’s onboard computing runs a 300M-parameter diffusion policy with about $100\leavevmode\nobreak\ \mathrm{ms}$ latency, enabling real-time operation.
+ToddlerBot’s onboard computing runs a 300300M parameter diffusion policy with about $100\leavevmode\nobreak\ \mathrm{ms}$ latency, enabling real-time operation.
 Results are shown in Figure 4.
-V-E ML-Compatibility: Skill Chaining
-To test the system’s ability to combine locomotion and manipulation skills, we test the wagon-pushing task (Figure 8), which requires the robot to first perform vision-based grasping and then walk forward.
-To do so, ToddlerBot first executes a diffusion policy to grasp the handle, while maintaining that pose, switched to the RL policy to push the wagon forward.
-To enable ToddlerBot to walk while maintaining its grip, we sample the end pose of the upper body from the grasping policy’s training data (60 demonstrations) during RL training.
-V-F Reproducibility: Hardware and Policies
-To demonstrate hardware reproducibility, we recruit a CS-major student with no prior hardware experience to assemble a second ToddlerBot using our manuals as shown in Figure 15 of the supplementary material and our open-source assembly videos.
-The student independently completes the assembly in three days, including the time required for 3D printing.
+ML-Compatibility: Skill Chaining We test ToddlerBot’s ability to combine loco-manipulation skills with wagon pushing.
+The results are shown in Appendix 8.15.
+To push the wagon, ToddlerBot first executes a diffusion policy to grasp the handle, while maintaining that pose, switches to the RL policy to walk forward.
+To enable walking while maintaining the grip, we sample the robot’s end pose from the grasping policy’s training data (60 demonstrations) during RL training.
+Reproducibility: Hardware and Policies To demonstrate hardware reproducibility, we recruit a CS-major student with no prior hardware experience to assemble a second ToddlerBot using our open-source assembly manual and videos.
+The student independently completes the assembly in three days, including the time for 3D printing.
+The open-source community has also reported successful reproductions of ToddlerBot, and most of them were completed in one week (Appendix 8.16).
 For policy reproducibility, we run the manipulation policy trained on data collected with one instance on the other instance, achieving the same success rate of 90% across 20 trials.
-We also successfully transfer the RL walking policy between both robots.
-To further showcase the equivalent performance of both ToddlerBot instances, we have them collaborate on a long-horizon room tidying task, as shown in Figure 9.
-VI Limitations and Future Work
+We also successfully transferred the RL walking policy between both robots.
+To further showcase the equivalent performance of both ToddlerBot instances, we have them collaborate on a long-horizon toy tidy-up session, as shown in Figure 5.
+6 Conclusions
+In conclusion, we demonstrate that ToddlerBot is ML-compatible, capable, and reproducible through a series of tests and loco-manipulation tasks.
+ToddlerBot extends beyond locomotion and supports full-body manipulation, character animation, human-robot interaction, and a range of ML applications, making it a versatile research platform.
+To bridge the sim-to-real gap, we develop a comprehensive sysID pipeline to create a high-fidelity digital twin, alongside a whole-body teleoperation system to collect loco-manipulation data.
+Being fully open-source, ToddlerBot empowers researchers to explore new research directions and fosters open collaboration across the community.
+7 Limitations and Future Work
 ToddlerBot’s performance in more agile tasks is constrained by the off-the-shelf motors’ max speed, max torque, and communication speed.
+Should there be stronger motor choices, the performance can still improve.
 Rather than achieving superhuman capabilities, ToddlerBot aligns more closely with average human performance in loco-manipulation tasks.
-Additionally, its scale limits interaction with human-sized objects, though this does not hinder research if appropriately sized objects are used.
-To overcome these limitations, we are developing customized communication boards and improving motor system identification to maximize performance.
-We also aim to improve sensing capabilities, including stereo vision for depth perception, additional IMUs for improved state estimation, and tactile sensors for richer feedback.
-VII Conclusions
-In conclusion, we demonstrate that ToddlerBot is ML-compatible, capable, and reproducible through a series of tests and loco-manipulation tasks.
-While humanoid research is often associated with locomotion, ToddlerBot extends beyond this to support full-body manipulation, character animation, human-robot interaction, and various ML applications, making it a versatile research platform.
-With ToddlerBot fully open-source, we hope to empower researchers to explore new directions in humanoid research and encourage open collaboration in the community.
+In addition, the current actuation model does not consider motor temperature and is less accurate when close to the performance limit.
+The miniature scale limits its interaction with human-sized objects, though this does not hinder research if appropriately sized objects are used.
+Compared to metal shells, 3D-printed parts are more likely to break after the event of impact.
+Despite 3D printing being fast, repairs still require time.
+To overcome these limitations, we plan to develop customized communication boards to increase control frequency and improve our actuation model to maximize usable envelope, especially when the motor heats up.
+We will refine ToddlerBot’s design to improve structural strength, as we believe that design has a greater impact than material choice.
+We also aim to improve sensing capabilities, including stereo vision for depth perception, additional IMUs for improved state estimation, and tactile sensors for richer manipulation feedback.
 Acknowledgments
@@
 Finally, we appreciate the helpful discussions from all members of TML and REALab.
-This work was supported by National Science Foundation NSF-FRR-2153854, NSF-2143601, NSF-2037101, Sloan Fellowship, Stanford Institute for Human-Centered Artificial Intelligence, and Stanford Wu Tsai Human Performance Alliance.
+This work was supported by National Science Foundation NSF-FRR-2153854, NSF-2143601, NSF-2037101, NSF-2132519, Sloan Fellowship, Stanford Institute for Human-Centered Artificial Intelligence, and Stanford Wu Tsai Human Performance Alliance.

### v4

@@
 With its data collection ability and anthropomorphic design, ToddlerBot is ideal for whole-body loco-manipulation research.
-Additionally, ToddlerBot’s compact size ($0.56\leavevmode\nobreak\ \mathrm{m},3.4\leavevmode\nobreak\ \mathrm{kg}$) ensures safe operation in real-world environments.
-Reproducibility is achieved with entirely 3D-printed, open-source design and off-the-shelf components, keeping the total cost under $6,000\leavevmode\nobreak\ \mathrm{USD}$.
+Additionally, ToddlerBot’s compact size ($0.56\penalty 10000\ \mathrm{m},3.4\penalty 10000\ \mathrm{kg}$) ensures safe operation in real-world environments.
+Reproducibility is achieved with entirely 3D-printed, open-source design and off-the-shelf components, keeping the total cost under $6,000\penalty 10000\ \mathrm{USD}$.
 This allows assembly and maintenance with basic technical expertise, as validated by successful independent replications of the system.
@@
 We design an intuitive teleoperation interface that allows simultaneous control of ToddlerBot’s upper and lower body to collect whole-body manipulation data and develop effective visuomotor policies.
-Additionally, ToddlerBot’s small size and weight ($0.56\leavevmode\nobreak\ \mathrm{m},3.4\leavevmode\nobreak\ \mathrm{kg}$) ensure safe and accessible operation in real-world environments.
+Additionally, ToddlerBot’s small size and weight ($0.56\penalty 10000\ \mathrm{m},3.4\penalty 10000\ \mathrm{kg}$) ensure safe and accessible operation in real-world environments.
 Beyond ML-compatibility, ToddlerBot is designed with a focus on capability and reproducibility.
@@
 Reproducibility is achieved through low-cost, open-source designs and readily accessible hardware components.
-ToddlerBot uses commercially available motors and is completely 3D-printed, with a total cost under $6,000\leavevmode\nobreak\ \mathrm{USD}$ ($90\%$ of the cost is for motors and computers).
+ToddlerBot uses commercially available motors and is completely 3D-printed, with a total cost under $6,000\penalty 10000\ \mathrm{USD}$ ($90\%$ of the cost is for motors and computers).
 We release digital twin, learning algorithms, hardware designs, and comprehensive tutorials to ensure that ToddlerBot can be built at home with basic knowledge of hardware and software, without requiring specialized equipment for manufacturing or repair.
@@
 For performance-critical components like motors and bearings, we limit ourselves to off-the-shelf parts.
-The total BOM cost of ToddlerBot is $6,000\leavevmode\nobreak\ \mathrm{USD}$, with 90% spent on the computer (NVIDIA Jetson) and motors (Robotis Dynamixel).
+The total BOM cost of ToddlerBot is $6,000\penalty 10000\ \mathrm{USD}$, with 90% spent on the computer (NVIDIA Jetson) and motors (Robotis Dynamixel).
 We detail our design considerations in Appendix 8.2.
@@
 Figure 3: Arm Span, Payload, and Trajectory Tracking.
-On the left, we show that with a torso dimension of $13\times 9\times 12\leavevmode\nobreak\ \mathrm{cm}^{3}$, ToddlerBot can grasp objects up to $27\times 24\times 31\leavevmode\nobreak\ \mathrm{cm}^{3}$, about 14 times the torso size.
-Additionally, ToddlerBot can lift weights up to $1484\leavevmode\nobreak\ \mathrm{g}$, which is 40% of its body weight ($3484\leavevmode\nobreak\ \mathrm{g}$).
+On the left, we show that with a torso dimension of $13\times 9\times 12\penalty 10000\ \mathrm{cm}^{3}$, ToddlerBot can grasp objects up to $27\times 24\times 31\penalty 10000\ \mathrm{cm}^{3}$, about 14 times the torso size.
+Additionally, ToddlerBot can lift weights up to $1484\penalty 10000\ \mathrm{g}$, which is 40% of its body weight ($3484\penalty 10000\ \mathrm{g}$).
 On the right, we present ten consecutive real-world rollouts of an RL walking policy tracking a square trajectory with a predefined velocity profile.
@@
 To evaluate ToddlerBot’s arm span, we teleoperate it to hug a large object using the compliant palm gripper while maintaining balance.
-We show that, ToddlerBot can grasp objects up to $27\times 24\times 31\leavevmode\nobreak\ \mathrm{cm}^{3}$, approximately 14 times its torso volume.
+We show that, ToddlerBot can grasp objects up to $27\times 24\times 31\penalty 10000\ \mathrm{cm}^{3}$, approximately 14 times its torso volume.
 The payload test assesses both the upper body’s lifting capacity and the lower body’s ability to maintain balance.
-ToddlerBot successfully lifts up to $1484\leavevmode\nobreak\ \mathrm{g}$, 40% of its total weight ($3484\leavevmode\nobreak\ \mathrm{g}$).
+ToddlerBot successfully lifts up to $1484\penalty 10000\ \mathrm{g}$, 40% of its total weight ($3484\penalty 10000\ \mathrm{g}$).
 To eliminate friction effects, we use a 3D-printed cup for the gripper to lock in securely.
@@
 However, the sim-to-real gap is notably smaller than the tracking gap, supporting a successful zero-shot transfer.
-Additionally, we report a position tracking error variance of $0.018\leavevmode\nobreak\ \mathrm{m}$, which demonstrates good repeatability (more details in Appendix 8.14).
+Additionally, we report a position tracking error variance of $0.018\penalty 10000\ \mathrm{m}$, which demonstrates good repeatability (more details in Appendix 8.14).
 ML-Compatibility: Vision-based Manipulation.
@@
 In the bimanual task, the torso rotating and releasing motions are open-loop, while in full-body manipulation, kneeling down is open-loop.
-ToddlerBot’s onboard computing runs a 300300M parameter diffusion policy with about $100\leavevmode\nobreak\ \mathrm{ms}$ latency, enabling real-time operation.
+ToddlerBot’s onboard computing runs a 300300M parameter diffusion policy with about $100\penalty 10000\ \mathrm{ms}$ latency, enabling real-time operation.
 Results are shown in Figure 4.